SubjectCover

Question 18 of 19 in Handling Data 5

Handling Data 5

4 marks · Hard

A box contains 3 orange balls and 2 purple balls. A ball is picked at random, not replaced, then a second ball is picked at random.

Calculate the probability that: (a) Both balls are orange (b) The two balls picked are different colours

(a)

Calculate the probability that both balls are orange.

[2 marks]
(b)

Calculate the probability that the two balls are different colours.

[2 marks]

Work this out on paper, then compare with the answer and worked solution below. With a free account you can submit written answers for marking.

Show answer

(a) 3/10

(b) 6/10 or 3/5

Common mistakes: Incorrectly using replacement (independent events); Errors in fractions or numerators/denominators

Show mark scheme and worked solution

Marking points

M1 (a) · 1 mark · Multiplies correct probabilities for first and second orange ball (3/5 × 2/4 or equivalent).

A1 (a) · 1 mark · Gives final answer 3/10 for both orange.

M2 (b) · 1 mark · Adds probabilities for orange then purple, and purple then orange, or finds 1 minus probability of both same. (follow-through)

A2 (b) · 1 mark · Gives final answer 6/10 or 3/5 for different colours. (follow-through)

Worked solution

(a) Probability of first ball orange: 3/5. Probability of second ball orange (after one orange removed): 2/4 = 1/2.

So, probability both orange = 3/10.

(b) Probability of first orange then second purple: (3/5) × (2/4) = 3/10 Probability of first purple then second orange: (2/5) × (3/4) = 6/20 = 3/10 Total probability = 3/10 + 3/10 = 6/10 = 3/5

Report a problem with this question
What's wrong?